Gordon Co., having recently issued a $1,000,000 bond, is committed to make annual sinking fund deposits...

  1. Gordon Co., having recently issued a $1,000,000 bond, is committed to make annual sinking fund deposits of $75,000.  The deposits are made on the last day of each year and yield a return of 6%.  Will the fund at the end of 10 years be sufficient to retire the bonds?  If not, what will the deficiency be?

  1. LakeFront Co. is considering investing in a new dock that will cost $280,000.  The company expects to use the dock for 5 years, after which it will be sold for $150,000 at that time.  LakeFront anticipates cash flows of $50,000 resulting from the new dock.  If the appropriate interest rate is 6%, would it be worth buying the dock?

  1. Assume that JJ Inc. has a contractual debt outstanding with two options for settlement:  (a) pay $3,500,000 today or (b) make annual payments of $400,000 for 15 years, with each payment due at the end of the year.  Assuming an effective rate of 8%, which would you recommend?
Answer #1

1: Using financial calculator

Input: PMT=-75000

N= 10

I/Y=6

Solve for FV as $988,559.62

The amount is not sufficient to retire the bonds.

Deficiency = 1000000- 988,559.62 = $11,440.38

2: Using financial calculator

Input:I/Y=6, N=5, PMT=-50000

Solve for PV as $210,618.19

PV of cashflows= $210,618.19

PV of sale amount = 150000/1.06^5 = $112,088.73

Total PV = $322,706.92

Hence it is worth investing in the stock since PV of cash inflows is more than the cost.

3: PV of annual payments = PV of annuity = Annuity*(1-1/(1+rate)^number of terms)/rate

= 400000*(1-1/(1+8%)^15)/8%

= 3423791.48

Option b is better since its present value is lower.